# Goal-guided Generative Prompt Injection Attack on Large Language Models

### v2

@@ Abstract
Current large language models (LLMs) provide a strong foundation for large-scale user-oriented natural language tasks. Numerous users can easily inject adversarial text or instructions through the user interface, thus causing LLM model security challenges. Although there is currently much research on prompt injection attacks, most of these black-box attacks use heuristic strategies. It is unclear how these heuristic strategies relate to the success rate of attacks and thus effectively improve model robustness. To solve this problem, we redefine the goal of the attack: to maximize the KL divergence between the conditional probabilities of the clean text and the adversarial text. Furthermore, we prove that maximizing the KL divergence is equivalent to maximizing the Mahalanobis distance between the embedded representation $x$ and $x^{\prime}$ of the clean text and the adversarial text when the conditional probability is a Gaussian distribution and gives a quantitative relationship on $x$ and $x^{\prime}$. Then we designed a simple and effective goal-guided generative prompt injection strategy (G2PIA) to find an injection text that satisfies specific constraints to achieve the optimal attack effect approximately. It is noteworthy that our attack method is a query-free black-box attack method with a low computational cost. Experimental results on seven LLM models and four datasets show the effectiveness of our attack method.
@@ I Introduction
In our work, we assume that the clean text representation $x$ and the adversarial text representation $x^{\prime}$ satisfy the conditional probability distribution $p(y|x)$ and $p(y|x^{\prime})$ respectively, and the goal of the black-box attack is to maximize the KL divergence $\textrm{KL}(p(y|x),p(y|x^{\prime}))$, then we prove that maximizing the KL divergence is equivalent to maximizing the Mahalanobis distance between $x$ and $x^{\prime}$ under the assumption of Gaussian distribution. Furthermore, we give the quantitative relationship between optimal attack text representation ${x^{\prime}}^{*}$ and $x$. Based on the above theoretical results, we designed a simple and effective prompt text injection method to search for attack texts that meet approximately optimal conditions.
Overall, our contributions are as follows: 1) We propose a new objective function based on KL divergence between two conditional probabilities for black-box attacks to maximize the success rate of black-box attacks; 2) We theoretically prove that under the assumption that the conditional probabilities are Gaussian distributions, the KL divergence maximization problem based on the posterior probability distributions of clean text and adversarial text, respectively, is equivalent to maximizing the Mahalanobis distance between clean text and adversarial text. 3) We propose a simple and effective injection attack strategy for generating adversarial injection texts, and the experimental results verify the effectiveness of our method. Note that our attack method is a query-free black-box attack method with low computational cost.
@@ III-B Analysis on Objective
To solve the problem (5) theoretically, we do the following processing: Replace the discrete Gaussian distribution with a continuous Gaussian distribution to facilitate the calculation of KL divergence although the output text on condition of input text still obeys discrete distribution in practical applications. Subsequently, we present the following theorem (see Appendix Section A and B for details).
@@ IV-A1 Victim Models
  • Llama-2. Llama-2 [26] is Meta AI’s next-generation open-source large language model. It is more capable than Llama 1 and outperforms other open-source language models on several external benchmarks, including reasoning, encoding, proficiency, and knowledge tests. Llama 2-7B, Llama 2-13B, and Llama 2-70B are transformer framework-based models.
@@ VI Conclusion
In our work, we propose a new goal-oriented generative prompt injection attack (G2PIA) method. To make the injected text mislead the large model as much as possible, we define a new objective function to maximize, which is the KL divergence value between the two posterior conditional probabilities before injection (clean text) and after injection (attack text). Furthermore, we proved that under the condition that the conditional probability follows the multivariate Gaussian distribution, maximizing the KL divergence value is equivalent to maximizing the Mahalanobis distance between clean text and adversarial text. Then, we establish the relationship between the optimal adversarial text and clean text. Based on the above conclusions, we design a simple and effective attack strategy with an assisted model to generate injected text that satisfies certain constraints, maximizing the KL divergence. Experimental results on multiple public datasets and popular LLMs demonstrate the effectiveness of our method.
@@ A-C Solution of Problem (24) in Two Dimensional Space
A-D Relationship between KL Divergence and Cos Similarity in 2-dimensional space
When the point $z$ moves on the ellipse, its length corresponds to the solution of the optimization problem Eqn. (24): when it takes the minimum value, the optimization problem takes the maximum value. It can be seen from Fig. 2 that when pointing $z$ moves from $A$ to $B$, $\cos(x^{\prime},x)$ gradually increases, but the length of $z$ first decreases and then increases. This process corresponds to that as the cos similarity increases, the KL divergence value first increases from a non-optimal value to the maximum value, then decreases. We experimentally observed the attack effect as shown in Fig. 6, verifying our conclusion.
@@ A-E Impact of Parameters $(\epsilon,\gamma)$ on Attack Effects
A-F Relationship between KL Divergence and Cos Similarity in Two-dimensional space
When point $z$ moves on the ellipse, its length corresponds to the solution of the optimization problem Eqn. (24): when it takes the minimum value, the optimization problem takes the maximum value. It can be seen from Fig. 2 that when pointing $z$ moves from $A$ to $B$, $\cos(x^{\prime},x)$ gradually increases, but the length of $z$ first decreases and then increases. This process corresponds to that as the cos similarity increases, the KL divergence value first increases from a non-optimal value to the maximum value, then decreases. We experimentally observed the attack effect as shown in Fig. 6, verifying our conclusion.

### v4

@@ Abstract
Current large language models (LLMs) provide a strong foundation for large-scale user-oriented natural language tasks. Numerous users can easily inject adversarial text or instructions through the user interface, thus causing LLM model security challenges. Although there is currently much research on prompt injection attacks, most of these black-box attacks use heuristic strategies. It is unclear how these heuristic strategies relate to the success rate of attacks and thus effectively improve model robustness. To solve this problem, we redefine the goal of the attack: to maximize the KL-divergence between the conditional probabilities of the clean text and the adversarial text. Furthermore, we prove that maximizing the KL-divergence is equivalent to maximizing the Mahalanobis distance between the embedded representation $x$ and $x^{\prime}$ of the clean text and the adversarial text when the conditional probability is a Gaussian distribution and gives a quantitative relationship on $x$ and $x^{\prime}$. Then we designed a simple and effective goal-guided generative prompt injection strategy (G2PIA) to find an injection text that satisfies specific constraints to achieve the optimal attack effect approximately. It is noteworthy that our attack method is a query-free black-box attack method with a low computational cost. Experimental results on seven LLM models and four datasets show the effectiveness of our attack method.
@@ I Introduction
In our work, we assume that the clean text representation $x$ and the adversarial text representation $x^{\prime}$ satisfy the conditional probability distribution $p(y|x)$ and $p(y|x^{\prime})$ respectively, and the goal of the black-box attack is to maximize the KL-divergence $\textrm{KL}(p(y|x),p(y|x^{\prime}))$, then we prove that maximizing the KL-divergence is equivalent to maximizing the Mahalanobis distance between $x$ and $x^{\prime}$ under the assumption of Gaussian distribution. Furthermore, we give the quantitative relationship between optimal attack text representation ${x^{\prime}}^{*}$ and $x$. Based on the above theoretical results, we designed a simple and effective prompt text injection method to search for attack texts that meet approximately optimal conditions.
Overall, our contributions are as follows: 1) We propose a new objective function based on KL-divergence between two conditional probabilities for black-box attacks to maximize the success rate of black-box attacks; 2) We theoretically prove that under the assumption that the conditional probabilities are Gaussian distributions, the KL-divergence maximization problem based on the posterior probability distributions of clean text and adversarial text, respectively, is equivalent to maximizing the Mahalanobis distance between clean text and adversarial text. 3) We propose a simple and effective injection attack strategy for generating adversarial injection texts, and the experimental results verify the effectiveness of our method. Note that our attack method is a query-free black-box attack method with low computational cost.
@@ III-B Analysis on Objective
To solve the problem (5) theoretically, we do the following processing: Replace the discrete Gaussian distribution with a continuous Gaussian distribution to facilitate the calculation of KL-divergence although the output text on condition of input text still obeys discrete distribution in practical applications. Subsequently, we present the following theorem (see Appendix Section A and B for details).
@@ IV-A1 Victim Models
  • Llama-2. Llama-2 [26] is Meta AI’s next-generation open-source large language model. It is more capable than Llama and outperforms other open-source language models on several external benchmarks, including reasoning, encoding, proficiency, and knowledge tests. Llama 2-7B, Llama 2-13B, and Llama 2-70B are transformer framework-based models.
@@ VI Conclusion
In our work, we propose a new goal-oriented generative prompt injection attack (G2PIA) method. To make the injected text mislead the large model as much as possible, we define a new objective function to maximize, which is the KL-divergence value between the two posterior conditional probabilities before injection (clean text) and after injection (attack text). Furthermore, we proved that under the condition that the conditional probability follows the multivariate Gaussian distribution, maximizing the KL-divergence value is equivalent to maximizing the Mahalanobis distance between clean text and adversarial text. Then, we establish the relationship between the optimal adversarial text and clean text. Based on the above conclusions, we design a simple and effective attack strategy with an assisted model to generate injected text that satisfies certain constraints, maximizing the KL-divergence. Experimental results on multiple public datasets and popular LLMs demonstrate the effectiveness of our method.
@@ A-C Solution of Problem (24) in Two Dimensional Space
A-D Relationship between KL-divergence and Cos Similarity in 2-dimensional space
When the point $z$ moves on the ellipse, its length corresponds to the solution of the optimization problem Eqn. (24): when it takes the minimum value, the optimization problem takes the maximum value. It can be seen from Fig. 2 that when pointing $z$ moves from $A$ to $B$, $\cos(x^{\prime},x)$ gradually increases, but the length of $z$ first decreases and then increases. This process corresponds to that as the cos similarity increases, the KL-divergence value first increases from a non-optimal value to the maximum value, then decreases. We experimentally observed the attack effect as shown in Fig. 6, verifying our conclusion.
@@ A-E Impact of Parameters $(\epsilon,\gamma)$ on Attack Effects
A-F Relationship between KL-divergence and Cos Similarity in Two-dimensional space
When point $z$ moves on the ellipse, its length corresponds to the solution of the optimization problem Eqn. (24): when it takes the minimum value, the optimization problem takes the maximum value. It can be seen from Fig. 2 that when pointing $z$ moves from $A$ to $B$, $\cos(x^{\prime},x)$ gradually increases, but the length of $z$ first decreases and then increases. This process corresponds to that as the cos similarity increases, the KL-divergence value first increases from a non-optimal value to the maximum value, then decreases. We experimentally observed the attack effect as shown in Fig. 6, verifying our conclusion.